# Robust Instance Optimal Phase-Only Compressed Sensing

### v2

@@ 3.2 Bounded Dense Noise
Next, we proceed to explore the robustness to noise and corruption. We consider the nonuniform robustness concerning the reconstruction of a fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$, which restricts our attention to robustness without being distructed by the concerns of uniformity or instance optimality. We will discuss in Section 4 the work needed to improve these forthcoming nonuniform robustness results to uniform ones with instance optimality.
We first consider pre-sign small dense noise $\bm{\tau}\in\mathbb{C}^{m}$ obeying $\|\bm{\tau}\|_{\infty}\leq\tau_{0}$, that is, observations given by $\breve{\mathbf{z}}=\mathbf{z}+\bm{\tau}$ [^5] We emphasize that the only constraint on $\bm{\tau}$ is a small enough max norm; under this constraint, it can be generated by an adversary having full knowledge of $(\bm{\Phi},\mathbf{x})$. This has been treated in [27, Sec. IV] and [10, Thm. 3]. We reproduce the result and proof here to demonstrate the two-step analysis for noisy PO-CS from [27]:
  • (1)
    Show the RIP of $\mathbf{A}_{\breve{\mathbf{z}}}$—based on the RIP of $\mathbf{A}_{\mathbf{z}}$ (Corollary 1) we only need to control the impact of noise;
  • (2)
    Estimate $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$, which is the noise level in the resulting linear compressed sensing problem, to indicate the choice of $\epsilon$ in (1.8).
@@
By the linearity of $\mathbf{A}_{\mathbf{u}}$ on $\mathbf{u}$ and $\breve{\mathbf{z}}=\mathbf{z}+\bm{\tau}$, we have $\mathbf{A}_{\breve{\mathbf{z}}}=\mathbf{A}_{\mathbf{z}}+\mathbf{A}_{\bm{\tau}}$. We proceed in several steps.
Show $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$: We first show that $\mathbf{A}_{\bm{\tau}}$ cannot have great affect on the RIP of $\mathbf{A}_{\mathbf{z}}$ under small enough $\tau_{0}$. To that end, we need to bound $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\tau}}\mathbf{u}\|_{2}$. By (1.7) and triangle inequality, we have
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\tau}}\mathbf{u}\|_{2}$ $\leq\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\bm{\tau}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\bm{\tau}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}$ (3.5)
$\leq\tau_{0}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\|_{1}}{\kappa m}+\tau_{0}\left(\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Re}\mathbf{u}\|_{2}}{\sqrt{m}}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Im}\mathbf{u}\|_{2}}{\sqrt{m}}\right).$
@@
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\|_{1}}{\kappa m}\leq 1+c_{1}$
@@
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Re}\mathbf{u}\|_{2}}{\sqrt{m}}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Im}\mathbf{u}\|_{2}}{\sqrt{m}}\leq 2+c_{3}$
for some $c_{3}$ that can be set sufficiently small, with probability at least $1-\exp(-s\log(\frac{en}{s}))$. Setting $c_{1}=c_{3}=\frac{1}{2}$, we obtain $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\bm{\tau}}\mathbf{u}\|_{2}\leq 4\tau_{0}\leq\frac{1}{9}$ with the promised probability. We now invoke Corollary 1 to ensure small enough $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}|\|\mathbf{A}_{\mathbf{z}}\mathbf{u}\|_{2}-1|$ when $m\geq C_{1}s\log(\frac{en}{s})$ for large enough $C_{1}$, with the desired probability. Taken collectively, we find that $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$.
Bound on $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$: Next, we bound
$\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\|\mathbf{A}_{\mathbf{z}}\mathbf{x}^{\star}+\mathbf{A}_{\bm{\tau}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\|\mathbf{A}_{\bm{\tau}}\mathbf{x}^{\star}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\|\mathbf{A}_{\bm{\tau}}\mathbf{x}\|_{2}$ (3.6)
for fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$ and the corresponding $\mathbf{x}^{\star}=\frac{\kappa m\cdot\mathbf{x}}{\|\bm{\Phi}\mathbf{x}\|_{1}}$. Again using (1.7) we have
$\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\|\mathbf{A}_{\tau}\mathbf{x}\|_{2}$ $\leq\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\left[\frac{|\Re(\bm{\tau}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\bm{\tau}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\right]$
$\leq\tau_{0}+\tau_{0}\cdot\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\cdot\frac{\|\bm{\Phi}\mathbf{x}\|_{2}}{\sqrt{m}}\leq\frac{5\tau_{0}}{2},$ (3.7)
where in the last inequality we use sub-Gaussian tail bounds to ensure sufficiently small $|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|$ (recall that $\|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1\|_{\psi_{2}}=O(1/\sqrt{m})$ due to (2.1)), and Bernstein’s inequality [49, Thm. 2.8.1] to ensure sufficiently small $|\frac{\|\bm{\Phi}\mathbf{x}\|_{2}^{2}}{m}-2|$, with the promised probability.
Now by Proposition 2.1 we have $\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}\leq 7\cdot\frac{5\tau_{0}}{2}=\frac{35}{2}\tau_{0}$. Then by (2.12) and the condition of $\|\mathbf{x}^{\star}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}$ being sufficiently close to $1$ we have
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}=\left\|\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}-\frac{\mathbf{x}^{\star}}{\|\mathbf{x}^{\star}\|_{2}}\right\|_{2}\leq\frac{2\|\hat{\mathbf{x}}-\mathbf{x}^{\star}\|_{2}}{\|\mathbf{x}^{\star}\|_{2}}\leq 36\tau_{0},$ (3.8)
as claimed. ∎
@@
We move to post-sign small dense noise. We again denote the dense noise by $\bm{\tau}$, but now the noisy observations are $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau})$. The robustness in this regime is less straightforward than the post-sign noise. The reason is that a small enough pre-sign perturbation $\tau$ can still greatly affect $\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}+\tau)$ if $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ is small. This makes the RIP of $\mathbf{A}_{\breve{\mathbf{z}}}$ less evident, for which we have to separately treat a small fraction of measurements with small $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ and the majority with $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ bounded away from 00. On the other hand, it comes a bit surprising that the error horizon remains at $O(\tau_{0})$, since some algebra with (2.12) finds $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=O(\tau_{0})$.
@@
We first transfer $\bm{\tau}$ to post-sign noise by writing $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\tilde{\bm{\tau}}$, where
@@
The entries of $\tilde{\bm{\tau}}$ may not be uniformly small, but we can establish a decomposition $\tilde{\bm{\tau}}=\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}$ where $\|\tilde{\bm{\tau}}_{1}\|_{\infty}$ is small and $\tilde{\bm{\tau}}_{2}$ is sparse. For the fixed $\mathbf{x}\in\Sigma^{n,*}_{s}$ and for some $\eta_{0}>0$ to be chosen, we have
@@
$\mathbbm{P}\big{(}|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m\big{)}\geq\mathbbm{P}\Big{(}{\rm Bin}(m,\sqrt{2/\pi}\eta_{0})\leq\eta_{0}m\Big{)}\geq 1-\exp(-c_{1}\eta_{0}m)$ (3.9)
@@
$[\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}+\tau_{i})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})]\mathbbm{1}(|\bm{\Phi}_{i}^{*}\mathbf{x}|>\eta_{0}),$
and hence by (2.11) we obtain $\|\tilde{\bm{\tau}}_{1}\|_{\infty}\leq\frac{2\tau_{0}}{\eta_{0}}$. The observations can now be expressed as $\breve{\mathbf{z}}=\mathbf{z}+\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}.$
Show $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$: We now want to show $\mathbf{A}_{\breve{\mathbf{z}}}=\mathbf{A}_{\mathbf{z}}+\mathbf{A}_{\tilde{\bm{\tau}}_{1}}+\mathbf{A}_{\tilde{\bm{\tau}}_{2}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$. Similarly to the proof of Theorem 3.3, we can use Corollary 1 to obtain $\mathbf{A}_{\mathbf{z}}\sim\operatorname{RIP}(\bm{\Sigma}^{n}_{2s},c_{2})$ with sufficiently small $c_{2}$, and all that remains is to show $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{1}}\mathbf{u}\|_{2}$ and $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{2}}\mathbf{u}\|_{2}$ are both small enough. Now let us fix $\eta_{0}$ to be some small enough positive constant. With the promised probability, the arguments from the proof of Theorem 3.3 imply
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{1}}\mathbf{u}\|_{2}\leq 4\|\tilde{\bm{\tau}}_{1}\|_{\infty}\leq\frac{8\tau_{0}}{\eta_{0}},$
which is small enough because of sufficiently small $\tau_{0}$ (and a specified $\eta_{0}>0$).
Next, we need to make sure $\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{2}}\mathbf{u}\|_{2}$ is sufficiently small. We let $\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\in\{0,1\}^{m}$ be the vector whose $1$’s indicate the support of $\tilde{\bm{\tau}}_{2}$. Starting as in (3.5) while proceeding with $\|\tilde{\bm{\tau}}_{2}\|_{0}\leq\eta_{0}m$, $\|\tilde{\bm{\tau}}_{2}\|_{\infty}\leq 2$, and Cauchy-Schwarz inequality, we obtain
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{2}}\mathbf{u}\|_{2}\leq\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\tilde{\bm{\tau}}_{2}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\tilde{\bm{\tau}}_{2}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}$
$\leq\|\tilde{\bm{\tau}}_{2}\|_{2}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\odot\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\|_{2}}{\kappa m}+\|\tilde{\bm{\tau}}_{2}\|_{\infty}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\odot\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\|_{2}}{\sqrt{m}}$
@@
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\tilde{\bm{\tau}}_{2}}\mathbf{u}\|_{2}=O\left(\sqrt{\frac{s\log(\frac{en}{s})}{m}}+\sqrt{\eta_{0}\log(\frac{e}{\eta_{0}})}\right),$ (3.12)
which is small enough due to the scaling of $m$ and sufficiently small $\eta_{0}$. Overall, we have arrived at the desired $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$.
Bound on $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}$: Next, as with (3.6), we bound the $\ell_{2}$ measurement error at $\mathbf{x}^{\star}$,
$\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\|\mathbf{A}_{\tilde{\bm{\tau}}}\mathbf{x}\|_{2}\leq(1+c_{4})\left[\frac{|\Re(\tilde{\bm{\tau}}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\tilde{\bm{\tau}}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\right],$ (3.13)
where $c_{4}>0$ can be small enough due to the sub-Gaussian concentration of $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}$ about $1$. Letting $\tilde{\tau}_{i}=\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}+\tau_{i})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})$ be the $i$-th entry of $\tilde{\bm{\tau}}$, we have

### v3

@@ 3.2 Bounded Dense Noise
Next, we study robustness to noise and corruption in a nonuniform setting. Specifically, we consider the reconstruction of a fixed sparse signal $\mathbf{x}\in\Sigma^{n,*}_{s},$ which allows us to decouple the desiderata of uniformity, instance optimality and robustness, and simply focus on the robustness for now. We discuss in Section 4 how these nonuniform results can be extended to uniform, instance-optimal guarantees.
For convenience, we specialize Lemma 3.1 to our setting.
Lemma 3.2 (Specialization of Lemma 3.1).
Under complex Gaussian $\bm{\Phi}$ and a fixed $\mathbf{x}\in\Sigma^{n,*}_{s},$ let $\breve{\mathbf{z}}$ be a noisy version of $\mathbf{z}=\operatorname{sign}(\bm{\Phi}\mathbf{x})$, and suppose that we compute $\mathbf{x}^{\sharp}=\frac{\hat{\mathbf{x}}}{\|\hat{\mathbf{x}}\|_{2}}$ with $\hat{\mathbf{x}}$ obtained by solving $\Delta(\mathbf{A}_{\breve{\mathbf{z}}};\mathbf{e}_{1},\varepsilon)$ in (1.8). If (3.3) and
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re((\breve{\mathbf{z}}-\mathbf{z})^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}((\breve{\mathbf{z}}-\mathbf{z})^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}<\frac{1}{9}.$ (3.7)
hold, then with probability at least $1-C\exp(-cs\log(\frac{en}{s}))$, we have
$\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{14\|\bm{\Phi}\mathbf{x}\|_{1}\varepsilon}{\kappa m}.$ (3.8)
Proof.
The statement follows from Lemma 3.1 after the validation of $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$. By $\mathbf{A}_{\breve{\mathbf{z}}}=\mathbf{A}_{\mathbf{z}}+\mathbf{A}_{\breve{\mathbf{z}}-\mathbf{z}}$, we have
$\|\mathbf{A}_{\mathbf{z}}\mathbf{u}\|_{2}-\|\mathbf{A}_{\breve{\mathbf{z}}-\mathbf{z}}\mathbf{u}\|_{2}\leq\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{u}\|_{2}\leq\|\mathbf{A}_{\mathbf{z}}\mathbf{u}\|_{2}+\|\mathbf{A}_{\breve{\mathbf{z}}-\mathbf{z}}\mathbf{u}\|_{2}\,,\,\forall\mathbf{u}\in\Sigma^{n,*}_{2s}$.
By the RIP of $\mathbf{A}_{\mathbf{z}}$ in Corollary 1,
$1-c^{\prime}\leq\|\mathbf{A}_{\mathbf{z}}\mathbf{u}\|_{2}\leq 1+c^{\prime}\,,\,\forall\mathbf{u}\in\Sigma^{n,*}_{2s}\,,\quad\text{where~}c^{\prime}:=\min\Big\{\sqrt{\frac{4}{3}}-\frac{10}{9},\frac{8}{9}-\sqrt{\frac{2}{3}}\Big\}$ (3.9)
holds with the promised probability. By (1.7), triangle inequality and (3.7), we obtain
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\|\mathbf{A}_{\breve{\mathbf{z}}-\mathbf{z}}\mathbf{u}\|_{2}\leq\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re((\breve{\mathbf{z}}-\mathbf{z})^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}((\breve{\mathbf{z}}-\mathbf{z})^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}<\frac{1}{9}.$
Combining the preceding three displays yields $\sqrt{\frac{2}{3}}\leq\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{u}\|_{2}\leq\sqrt{\frac{4}{3}}\,,\,\forall\mathbf{u}\in\Sigma^{n,*}_{2s},$ which is exactly $\mathbf{A}_{\breve{\mathbf{z}}}\sim\operatorname{RIP}(\Sigma^{n}_{2s},\frac{1}{3})$. ∎
We first consider post-sign small dense noise $\bm{\tau}\in\mathbb{C}^{m}$ obeying $\|\bm{\tau}\|_{\infty}\leq\tau_{0}$, that is, observations given by $\breve{\mathbf{z}}=\mathbf{z}+\bm{\tau}$. [^6] We emphasize that the only constraint on $\bm{\tau}$ is a small enough max norm; under this constraint, it can be generated by an adversary having full knowledge of $(\bm{\Phi},\mathbf{x})$. This has been treated in [27, Sec. IV] and [10, Thm. 3]. We reproduce the result here as an illustration of the applicability of Lemma 3.2.
@@
By Lemma 3.2, we only need to establish (3.7) and (3.3) with $\varepsilon=\frac{5}{2}\tau_{0}$ with the promised probability: once these two conditions hold true, then by Lemma 3.2 we have $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{\|\bm{\Phi}\mathbf{x}\|_{1}35\tau_{0}}{\kappa m}$, and further using $\mathbb{P}(\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}\leq\frac{36}{35})\geq 1-\exp(-cm)$ yields the claim (recall that $\|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1\|_{\psi_{2}}=O(1/\sqrt{m})$ due to (2.1)).
Establish (3.7): Note that $\breve{\mathbf{z}}-\mathbf{z}=\bm{\tau}$ satisfies $\|\bm{\tau}\|_{\infty}\leq\tau_{0}\leq\frac{1}{36}$, we use triangle inequality to obtain
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\bm{\tau}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\bm{\tau}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}$
$\leq\tau_{0}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\|_{1}}{\kappa m}+\tau_{0}\left(\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Re}\mathbf{u}\|_{2}}{\sqrt{m}}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Im}\mathbf{u}\|_{2}}{\sqrt{m}}\right).$ (3.10)
@@
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\|_{1}}{\kappa m}\leq 1+c_{1}$ (3.11)
@@
$\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Re}\mathbf{u}\|_{2}}{\sqrt{m}}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}^{\Im}\mathbf{u}\|_{2}}{\sqrt{m}}\leq 2+c_{3}$ (3.12)
for some $c_{3}$ that can be set sufficiently small, with probability at least $1-\exp(-s\log(\frac{en}{s}))$. Setting $c_{1}=c_{3}=\frac{1}{2}$, we obtain (3.7) with the promised probability.
Establish (3.3) with $\varepsilon=\frac{5}{2}\tau_{0}$: This can be seen by
$\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\left[\frac{|\Re(\bm{\tau}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\bm{\tau}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\right]\leq\tau_{0}+\tau_{0}\cdot\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\cdot\frac{\|\bm{\Phi}\mathbf{x}\|_{2}}{\sqrt{m}}\leq\frac{5\tau_{0}}{2},$ (3.13)
where in the first inequality we use $|\Re(\bm{\tau}^{*}\bm{\Phi}\mathbf{x})|\leq\tau_{0}\|\bm{\Phi}\mathbf{x}\|_{1}$ and $\|\Im(\operatorname{diag}(\bm{\tau}^{*})\bm{\Phi}\mathbf{x})\|_{2}\leq\tau_{0}\|\bm{\Phi}\mathbf{x}\|_{2}$, in the second inequality we use $\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\cdot\frac{\|\bm{\Phi}\mathbf{x}\|_{2}}{\sqrt{m}}<\frac{3}{2}$ which holds because sub-Gaussian tail bound in (3.2) guarantees sufficiently small $|\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}-1|$, and Bernstein’s inequality [50, Thm. 2.8.1] yields $\mathbb{P}(|\frac{\|\bm{\Phi}\mathbf{x}\|_{2}^{2}}{m}-2|\geq t)\leq 2\exp(-c_{4}mt^{2})$ for any $t\in(0,1)$ and thus guarantees sufficiently small $|\frac{\|\bm{\Phi}\mathbf{x}\|_{2}^{2}}{m}-2|$ with the promised probability. ∎
@@
We move to pre-sign small dense noise. We again denote the dense noise by $\bm{\tau}$, but now the noisy observations are $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x}+\bm{\tau})$. The robustness in this regime is less straightforward than the post-sign noise. The reason is that a small enough pre-sign perturbation $\tau$ can still greatly affect $\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}+\tau)$ if $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ is small. This makes the RIP of $\mathbf{A}_{\breve{\mathbf{z}}}$ less evident, for which we have to separately treat a small fraction of measurements with small $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ and the majority with $|\bm{\Phi}_{i}^{*}\mathbf{x}|$ bounded away from 0. On the other hand, it comes a bit surprising that the error horizon remains at $O(\tau_{0})$, since some algebra with (2.12) finds $\|\mathbf{A}_{\breve{\mathbf{z}}}\mathbf{x}^{\star}-\mathbf{e}_{1}\|_{2}=O(\tau_{0})$.
@@
In light of Lemma 3.2, we only need to establish (3.7) and (3.3) with $\varepsilon=4\tau_{0}$ with the promised probability, then by Lemma 3.2 we have $\|\mathbf{x}^{\sharp}-\mathbf{x}\|_{2}\leq\frac{\|\bm{\Phi}\mathbf{x}\|_{1}\cdot 56\tau_{0}}{\kappa m}$, which leads to the claim on the event $\frac{\|\bm{\Phi}\mathbf{x}\|_{1}}{\kappa m}\leq\frac{57}{56}$ that holds with probability at least $1-2\exp(-cm)$.
To get started, we first transfer $\bm{\tau}$ to post-sign noise by writing $\breve{\mathbf{z}}=\operatorname{sign}(\bm{\Phi}\mathbf{x})+\tilde{\bm{\tau}}$, where
@@
The entries of $\tilde{\bm{\tau}}$ may not be uniformly small, but we can establish a decomposition $\tilde{\bm{\tau}}=\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}$ where $\|\tilde{\bm{\tau}}_{1}\|_{\infty}$ is small and $\tilde{\bm{\tau}}_{2}$ is sparse. We let $\eta_{0}:=144c_{0}$. For the fixed $\mathbf{x}\in\Sigma^{n,*}_{s},$ we have
@@
$\mathbbm{P}\big(|\mathcal{J}_{\mathbf{x},\eta_{0}}|\leq\eta_{0}m\big)\geq\mathbbm{P}\Big({\rm Bin}(m,\sqrt{2/\pi}\cdot\eta_{0})\leq\eta_{0}m\Big)\geq 1-\exp(-c_{1}\eta_{0}m)$ (3.14)
@@
$\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}+\tau_{i})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})]\mathbbm{1}(|\bm{\Phi}_{i}^{*}\mathbf{x}|>\eta_{0}),$
and hence by (2.11) we obtain $\|\tilde{\bm{\tau}}_{1}\|_{\infty}\leq\frac{2\tau_{0}}{\eta_{0}}$. The observations can now be expressed as
$\breve{\mathbf{z}}=\mathbf{z}+\tilde{\bm{\tau}}_{1}+\tilde{\bm{\tau}}_{2}.$ (3.15)
Establish (3.7): By (3.15) and triangle inequality, to guarantee (3.7), it is sufficient to establish
$T_{1}:=\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\tilde{\bm{\tau}}_{1}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\tilde{\bm{\tau}}_{1}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}\leq\frac{1}{18},$ (3.16)
$T_{2}:=\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{|\Re(\tilde{\bm{\tau}}_{2}^{*}\bm{\Phi}\mathbf{u})|}{\kappa m}+\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\Im(\operatorname{diag}(\tilde{\bm{\tau}}_{2}^{*})\bm{\Phi}\mathbf{u})\|_{2}}{\sqrt{m}}\leq\frac{1}{18}.$ (3.17)
[…]
$T_{2}$ $\leq\|\tilde{\bm{\tau}}_{2}\|_{2}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\odot\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\|_{2}}{\kappa m}+\|\tilde{\bm{\tau}}_{2}\|_{\infty}\sup_{\mathbf{u}\in\Sigma^{n,*}_{2s}}\frac{\|\bm{\Phi}\mathbf{u}\odot\mathbf{1}_{\operatorname{supp}(\tilde{\bm{\tau}}_{2})}\|_{2}}{\sqrt{m}}$
@@
$T_{2}=O\left(\sqrt{\frac{s\log(\frac{en}{s})}{m}}+\sqrt{\eta_{0}\log(\frac{e}{\eta_{0}})}\right),$ (3.20)
which is small enough due to the scaling of $m$, $\eta_{0}=144c_{0}$ and the small enough $c_{0}$.
Establish (3.3) with $\varepsilon=4\tau_{0}$: Next, we seek to show
$\frac{\kappa m}{\|\bm{\Phi}\mathbf{x}\|_{1}}\Big[\frac{|\Re(\tilde{\bm{\tau}}^{*}\bm{\Phi}\mathbf{x})|}{\kappa m}+\frac{\|\Im(\operatorname{diag}(\breve{\bm{\tau}}^{*})\bm{\Phi}\mathbf{x})\|_{2}}{\sqrt{m}}\Big]\leq 4\tau_{0}.$ (3.21)
Letting $\tilde{\tau}_{i}=\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x}+\tau_{i})-\operatorname{sign}(\bm{\Phi}_{i}^{*}\mathbf{x})$ be the $i$-th entry of $\tilde{\bm{\tau}}$, we have